1. Mukecha manufacturing company uses two direct cost categories direct materials and direct labor. Standards were developed for each direct cost category and used for exercising control over these costs. The following are standards set by the company on January 1,2015, the first fiscal period of the company:

Direct material use                                                                                                        4kg per unit of output

Direct materials price per kg                                                                                       $60

Direct labor                                                                                                                    2 hours per unit of output

Direct labor rate per hour                                                                                           $40

               Actual data:

                           Actual units of output                                                                                                   22,000

                           Direct materials purchased($62/kg)                                                                           90,000kgs

                           Direct materials used during the year                                                                        80,000kgs

                           Direct labor hours used during the period                                                                 40,000hours

                           Actual direct labor rate per hour                                                                                 $36

Instruction:

  1. For direct materials, compute:
  1. Price variance
  2. Efficiency variance
  3. Flexible budget variance
  1. For direct labor, compute
  1. Rate variance
  2. Efficiency variance
  3. Flexible budget variance
Answer #1
DIRECT MATERIALS:
a) Material price variance = (Actual price-Standard price)*Actual quantity purchased =
= (62-60)*90000 = $       1,80,000 Unfavorable
b) Material efficiency variance = (Actual quantity used-Standard quantity)*Standard price = (80000-22000*4)*60 = $       4,80,000 Favorable
c) Flexible budget variance = 480000-180000 = $       3,00,000 Favorable
DIRECT LABOR:
a) Direct labor rate variance = (Actual rate-Standard rate)*Actual DLH = (36.00-40.00)*40000 = $       1,60,000 Favorable
Direct labor efficiency variance = (Actual DLH-Standard DLH)*Standard rate = (40000-22000*2)*40.00 = $       1,60,000 Favorable
flexible budget variance = 160000+160000 = $       3,20,000 Favorable